I need to know if this problem is identity or contradiction.
4 + 4 (n + 2) = 3n - 2 (n - 5)

Answer by tommyt3rd(5050)

4+4(n+2)=3n-2(n-5)
4+4n+8=3n-2n+10
4n+12=n+10
3n=-2
n=-2/3
neither, it is conditional